# custom.bib

### v1

@@ Abstract
Although existing model editing methods perform well in recalling exact edit facts, they often struggle in complex scenarios that require deeper semantic understanding rather than mere knowledge regurgitation. Leveraging the strong contextual reasoning abilities of large language models (LLMs), in-context learning (ICL) becomes a promising editing method by comprehending edit information through context encoding. However, this method is constrained by the limited context window of LLMs, leading to degraded performance and efficiency as the number of edits increases. To overcome this limitation, we propose InComeS [^1] https://github.com/Syon-Li/InComeS, a flexible framework that enhances LLMs’ ability to process editing contexts through explicit compression and selection mechanisms. Specifically, InComeS compresses each editing context into the key-value (KV) cache of a special gist token, enabling efficient handling of multiple edits without being restricted by the model’s context window. Furthermore, specialized cross-attention modules are added to dynamically select the most relevant information from the gist pools, enabling adaptive and effective utilization of edit information. We conduct experiments on diverse model editing benchmarks with various editing formats, and the results demonstrate the effectiveness and efficiency of our method. Corresponding author.
@@ 1 Introduction
Model editing, also known as knowledge editing, has seen rapid progress in recent years [alphaedit, coachhook, wise, knowedit]. Its primary goal is to precisely integrate updated knowledge into a model, enabling targeted behavioral modifications while maintaining performance on unrelated tasks. Existing techniques have demonstrated strong performance in accurately recalling edited facts [Yao, knowedit, edit-overfit]. However, they often struggle in more complex editing scenarios, such as multi-hop editing composition [mquake, edit-overfit], natural language editing [dune], and editing tasks that require reasoning and generalization [ripple-effects, knowedit]. Moreover, recent studies [edit-overfit] show that previous editing methods are prone to overfitting: they may assign excessively high probabilities to edited targets, which can distort the model’s responses to more complex or nuanced queries.
@@
To address these challenges, we introduce InComeS (Integrating Compression and Selection Mechanisms), a novel framework for efficient and scalable model editing. InComeS adopts context compression techniques to condense the representation of each edit into the KV cache of special gist tokens, which can be cached and reused for computational efficiency. While gisting [gisting] was originally developed to compress single-input prompts, we extend this approach to handle multiple edits by further introducing a specialized selection mechanism. Specifically, we augment the model with cross-attention modules that allow each input token to attend to the compressed gist representations of edits, enabling fine-grained and adaptive selection of the most relevant information. Since each edit is compressed in parallel, our framework overcomes the limitations imposed by the context window, and the specialized selection modules can be learned to enhance retrieval accuracy.
@@ 2 Preliminary
Model editing [Yao, mend] aims to adjust a base model $\psi$ to a post-edited model $\psi^{\prime}$ according to a set of editing information $\mathcal{T}=\{t_{1},\dots,t_{n}\}$:
@@ 3 Method
To enable accurate and efficient batch editing, we propose InComeS (Fig. 1), an ICL-based approach that integrates both compression and selection mechanisms into the LMs. First, we adopt gist-based edit compression, condensing each editing information into the KV cache representations of one special (gist) token. Furthermore, we introduce parallel-context cross-attention modules that allow ordinary tokens to attend to these compressed gist representations. These modules serve as soft selectors to dynamically identify the most relevant information for the current input. This strategy can effectively mitigate the limitations imposed by context window sizes and enhance the model’s ability to precisely capture editing information.
@@ 3.1 Edit Compression
We adopt the concept of gisting [gisting], which is originally developed to compress input prompts into the representations of an extra, specially inserted token (the gist token). The condensed gist activation serves the same function as the original prompt and can be cached for later reuse, thereby improving computational and memory efficiency. While the original work primarily focuses on instruction tuning, we extend this idea to edit compression.
@@ 3.3 Training
In addition to token reweighting, we also adopt knowledge distillation [distillation] to transfer the teacher model’s knowledge about the edit information into the target model. Specifically, we apply the KL divergence to align the output distributions of the gist-contextualized student model with those of the edit-contextualized teacher model:
@@
Here, $g_{1},\dots,g_{n}$ denote the cached gists for all the edits. We apply the token reweighting only to the vanilla cross-entropy term in our final loss, since we found that it would degrade effective learning of “attend-to-nothing” behavior if combined with the KL part.
@@ 4.1 Experiment setting
To verify the effectiveness of our method in complex editing scenarios, we conduct experiments on five popular datasets in model editing: the dataset for multi-hop editing MQuAKE [mquake], the natural language editing dataset DUNE [dune], the extended version of ZsRE [Yao, knowedit], which adds a portability test set to the original ZsRE [zsre], and the dataset containing ripple effect samples $\text{WikiData}_{counterfact}$ [ripple-effects, knowedit]. We report edit success rate and portability for the extend-ZsRE and $\text{WikiData}_{counterfact}$, the results for 2, 3, and 4 edits for MQuAKE [mquake] and new information, scientific reasoning, and debiasing for DUNE [dune]. More details about the datasets and evaluation metrics can be found in the Appendix.
@@
For baselines, we select representative methods demonstrated to be powerful in relevant surveys [Yao, knowedit]. For methods that directly edit the model weights, we include ROME [rome], R-ROME [r-rome], and MEMIT [memit]; for methods that adopt explicit external memory, we include SERAC [serac], and IKE [ike]; for methods that train additional meta-model or use implicit external memory (stores activations or nerons, etc), we adopt MEND [mend], GRACE [grace] and KN [kn]. We also include the traditional but powerful method, like fine-tuning, LoRA [LoRA], and ICL, which directly concatenates all the edits as the prefix context. We choose two representative open-source models for evaluation: Llama-3.2-1B  [^4] https://huggingface.co/meta-llama/Llama-3.2-1B and Qwen2.5-7B  [^5] https://huggingface.co/Qwen/Qwen2.5-7B [qwen2]. We apply cross-attention operations only for the second half of the model’s layers since we found that the gist KV cache from the first half is not informative enough to allow effective edit selection. For the calculation of cross-attention during inference, we adopt a temperature of $T=0.45$ to the logits before softmax, which is found to be helpful for effective editing. Unless otherwise specified, we adopt an edit batch size of 100 for batch editing. More details on the baseline implementation can be found in the Appendix.
@@ 4.2 Main results
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits 2-edits 3-edits 4-edits Base Llama-3.2-1B 41.79 43.51 31.58 41.79 43.51 31.58 FT-M 55.32 56.59 52.22 51.89 50.15 44.1 LoRA [LoRA] 67.63 68.84 60.21 50.29 47.88 47.15 ROME [rome] 1.97 7.77 0.55 - - - R-ROME [r-rome] 2.72 4.29 8.53 - - - MEMIT [memit] 40.21 39.54 22.34 29.71 28.87 15.7 EMMET [emmet] 5.4 6.61 4.74 11.49 17.16 15.89 MEND [mend] 39.88 35.23 33.31 35.55 30.29 29.89 GRACE [grace] 8.45 7.13 5.24 2.03 2.68 2.06 SERAC [serac] 41.84 43.51 31.58 41.84 43.45 31.74 ICL 59.23 59.00 51.63 50.06 49.86 42.37 InComeS $\textbf{71.19}_{20.19\%\uparrow}$ $\textbf{72.17}_{22.32\%\uparrow}$ $\textbf{72.62}_{40.65\%\uparrow}$ $\textbf{53.93}_{7.73\%\uparrow}$ $\textbf{52.79}_{5.88\%\uparrow}$ $\textbf{52.73}_{24.45\%\uparrow}$ Base Qwen2.5-7B 44.08 44.14 30.62 44.08 44.14 30.62 FT-M 69.89 74.88 74.96 50.31 49.04 50.39 LoRA [LoRA] 36.95 34.28 29.02 16.41 24.22 23.12 ROME [rome] 9.67 7.33 8.93 - - - R-ROME [r-rome] 10.73 6.68 3.62 - - - MEMIT [memit] 44.14 46.05 32.15 43.94 46.10 30.55 EMMET [emmet] 26.10 38.38 26.36 40.83 45.02 33.89 GRACE [grace] 15.78 13.45 13.23 5.43 7.88 4.23 SERAC [serac] 55.56 59.23 53.67 42.34 40.33 39.39 MEND [mend] 34.23 45.34 30.25 39.88 35.45 34.21 ICL 69.76 76.91 74.54 53.53 50.54 44.77 InComeS $66.46_{4.73\%\downarrow}$ $71.24_{7.37\%\downarrow}$ $\textbf{76.54}_{2.68\%\uparrow}$ $\textbf{55.13}_{3.00\%\uparrow}$ $\textbf{53.48}_{5.82\%\uparrow}$ $\textbf{47.91}_{7.01\%\uparrow}$
Table 1: Results on MQuAKE [mquake]. The difference between InComeS and ICL is marked.
@@
One of our method’s advantages is its flexibility in handling a variety of editing contexts with different formats. Unlike many traditional editing methods like ROME [rome] and MEMIT [memit], which require the input to follow the triplet-like fact statement format, InComeS can take edits in free-text forms without explicitly labeled subjects and objects. To verify our method’s capability for such scenarios, we adopt the DUNE dataset, which includes natural-language form edits, and the results are shown in Table 2. Following the original paper of DUNE [dune], we include fine-tuning, LoRA, SERAC [serac], and ICL as our baselines. The result confirms our method’s capability to handle natural language edits. Interestingly, the raw model itself is a strong baseline in the batch editing scenario, which may demonstrate the fast-evolving model capabilities over the years.
Method Model Single Editing Batch Editing New info Scientific R. Debiasing New info Scientific R. Debiasing Base Llama-3.2-1B 56.85 55.87 32.73 56.85 55.87 32.73 FT-M 57.43 53.34 35.73 57.07 53.45 33.43 LoRA 53.65 50.87 36.73 56.77 54.66 35.83 SERAC [serac] 52.67 48.96 46.73 50.76 47.32 36.52 ICL 58.67 55.84 56.65 56.94 55.68 39.46 InComeS $\textbf{60.00}_{2.27\%\uparrow}$ $\textbf{58.17}_{4.17\%\uparrow}$ $54.61_{3.60\%\downarrow}$ $\textbf{57.76}_{1.44\%\uparrow}$ $\textbf{56.46}_{1.40\%\uparrow}$ $\textbf{46.14}_{16.93\%\uparrow}$ Base Qwen2.5-7B 63.44 66.03 36.95 63.44 66.03 36.95 FT-M 64.83 67.77 48.73 64.27 64.56 41.51 LoRA 63.85 63.22 39.56 62.58 65.33 43.73 SERAC [serac] 64.45 63.57 33.38 56.78 58.97 31.24 ICL 65.81 65.34 35.25 66.29 66.24 40.73 InComeS $\textbf{66.83}_{1.55\%\uparrow}$ $\textbf{68.02}_{4.10\%\uparrow}$ $\textbf{62.59}_{77.56\%\uparrow}$ $65.61_{1.03\%\downarrow}$ $\textbf{67.82}_{2.39\%\uparrow}$ $\textbf{56.69}_{39.18\%\uparrow}$
Table 2: Results on DUNE [dune]. The difference between InComeS and ICL is marked.
@@
We further evaluate our method on two popular editing datasets that require reasoning abilities: $Wiki_{counterfact}$ [ripple-effects, knowedit] and the extended ZsRE [Yao, knowedit]. Table 3 shows the results. [^6] We only present methods deemed representative and powerful on these two datasets, detailed version can be found in Appendix. Our primary focus is on portability, as it serves as the most representative metric for assessing a model’s comprehensive understanding of the editing information. Overall, our method achieves performance comparable to ICL and consistently outperforms other baselines. As expected, traditional editing methods such as fine-tuning and ROME exhibit high edit success rates; however, their portability scores generally lag behind the top-performing methods, highlighting a common limitation of these approaches. In contrast, ICL-based approaches that leverage the in-context learning capabilities of LLMs demonstrate superior performance in complex editing scenarios that require reasoning, owing to the enhanced context understanding of LLMs.
Method Model $\text{WikiData}_{counterfact}$ ZsRE-extended Edit Success Portability Edit Success Portability Base Llama-3.2-1B 21.28 / 21.28 19.73 / 19.73 30.06 / 30.06 40.17 / 40.17 FT-M 97.02 / 94.58 53.43 / 47.51 99.81 / 95.94 62.80 / 54.84 LoRA [LoRA] 98.91 / 82.61 52.87 / 43.84 99.86 / 93.18 57.43 / 44.85 ROME [rome] 94.33 / 44-33 40.44 / 44-33 95.41 / 44-33 46.04 / 44-33 MEMIT [memit] 44-33 / 66.94 44-33 / 23.51 44-33 / 58.79 44-33 / 25.68 MEND [mend] 44-33 / 26.66 44-33 / 21.06 34-33 / 43.33 34-33 / 30.77 GRACE [grace] 33.27 / 25.06 14.33 / 10.51 32.00 / 24.44 12.73 / 10.91 IKE [ike] 61.70 / 34-33 45.55 / 34-33 59.15 / 43-33 57.39 / 43-33 SERAC [serac] 89.56 / 78.32 60.56 / 40.45 92.69 / 89.61 66.59 / 51.61 ICL 93.31 / 82.95 65.81 / 49.75 68.86 / 60.84 62.19 / 55.58 InComeS 91.16 / 76.81 65.15 / 45.66 97.22 / 87.09 70.70 / 52.23 Base Qwen2.5-7B 22.35 / 22.35 21.46 / 21.46 36.21 / 36.21 43.86 / 43.86 FT-M 98.93 / 90.18 49.39 / 43.13 99.51 / 92.60 50.04 / 46.41 LoRA [LoRA] 77.31 / 72.22 37.04 / 31.65 86.88 / 77.78 28.61 / 24.13 ROME [rome] 92.69 / 34-33 40.25 / 34-33 97.86 / 34-33 50.43 / 34-33 MEMIT [memit] 34-33 / 91.16 34-33 / 39.85 34-33 / 93.28 34-33 / 49.97 MEND [mend] 34-33 / 35.13 34-33 / 15.29 34-33 / 50.91 34-33 / 38.83 GRACE [grace] 31.34 / 33.77 25.60 / 18.55 33.27 / 26.79 14.35 / 11.25 IKE [ike] 96.40 / 34-33 75.33 / 34-33 99.75 / 34-33 83.17 / 34-33 SERAC [serac] 91.79 / 80.68 51.12 / 41.26 91.12 / 82.56 62.41 / 52.63 ICL 90.24 / 85.28 66.99 / 51.66 71.75 / 71.57 66.10 / 64.57 InComeS 90.96 / 71.44 66.69 / 47.93 97.95 / 91.29 75.63 / 61.22
Table 3: Results for $WikiData_{counterfact}$ [ripple-effects, knowedit] and ZsRE-extended [Yao, knowedit]. The data format of each cell is in "single/batch editing results". "-" means the methods are not designed for the corresponding settings. The best two statistics are marked.
@@
We further provide a scaling-up analysis to illustrate our method’s ability to generalize to larger numbers of edits, which is the main motivation of our modification over the ICL baseline. For this analysis, we use the COUNTERFACT dataset [memit], as it provides a sufficient number of editing instances. We vary the number of edits from 100 to 1000, resulting in total token counts ranging from approximately 1.2K to 12K. The results are shown in Figure 3, which shows that InComeS consistently outperforms ICL, though the base models have already been pretrained over long contexts [qwen2]. This finding suggests that the vanilla attention mechanism alone is insufficient to effectively comprehend and precisely select the required information from the context in complex editing scenarios. In contrast, our method demonstrates greater potential for handling large-scale edits through the unified compression and selection mechanism.

### v2

# InComeS: Integrating Compression and Selection Mechanisms into LLMs for Efficient Model Editing

@@ Abstract
Although existing model editing methods perform well in recalling exact edit facts, they often struggle in complex scenarios that require deeper semantic understanding rather than mere knowledge regurgitation. Leveraging the strong contextual reasoning abilities of large language models (LLMs), in-context learning (ICL) becomes a promising editing method by comprehending edit information through context encoding. However, this method is constrained by the limited context window of LLMs, leading to degraded performance and efficiency as the number of edits increases. To overcome this limitation, we propose InComeS, a flexible framework that enhances LLMs’ ability to process editing contexts through explicit compression and selection mechanisms. Specifically, InComeS compresses each editing context into the key-value (KV) cache of a special gist token, enabling efficient handling of multiple edits without being restricted by the model’s context window. Furthermore, specialized cross-attention modules are added to dynamically select the most relevant information from the gist pools, enabling adaptive and effective utilization of edit information. We conduct experiments on diverse model editing benchmarks with various editing formats, and the results demonstrate the effectiveness and efficiency of our method.
@@ 1 Introduction
Model editing, also known as knowledge editing, has seen rapid progress in recent years [8, 21, 40, 46]. Its primary goal is to precisely integrate updated knowledge into a model, enabling targeted behavioral modifications while maintaining performance on unrelated tasks. Existing techniques have demonstrated strong performance in accurately recalling edited facts [43, 46, 45]. However, they often struggle in more complex editing scenarios, such as multi-hop editing composition [49, 45], natural language editing [1], and editing tasks that require reasoning and generalization [2, 46]. Moreover, recent studies [45] show that previous editing methods are prone to overfitting: they may assign excessively high probabilities to edited targets, which can distort the model’s responses to more complex or nuanced queries.
@@
To address these challenges, we introduce InComeS  [^1] https://github.com/Syon-Li/InComeS (Integrating Compression and Selection Mechanisms), a novel framework for efficient and scalable model editing. InComeS adopts context compression techniques to condense the representation of each edit into the KV cache of special gist tokens, which can be cached and reused for computational efficiency. While gisting [30] was originally developed to compress single-input prompts, we extend this approach to handle multiple edits by further introducing a specialized selection mechanism. Specifically, we augment the model with cross-attention modules that allow each input token to attend to the compressed gist representations of edits, enabling fine-grained and adaptive selection of the most relevant information. Since each edit is compressed in parallel, our framework overcomes the limitations imposed by the context window, and the specialized selection modules can be learned to enhance retrieval accuracy.
@@ 2 Preliminary
Model editing [43, 28] aims to adjust a base model $\psi$ to a post-edited model $\psi^{\prime}$ according to a set of editing information $\mathcal{T}=\{t_{1},\dots,t_{n}\}$:
@@ 3 Method
To enable accurate and efficient batch editing, we propose InComeS an ICL-based approach that integrates both compression and selection mechanisms into the LMs. First, we adopt gist-based edit compression, condensing each editing information into the KV cache representations of one special (gist) token. Furthermore, we introduce parallel-context cross-attention modules that allow ordinary tokens to attend to these compressed gist representations. These modules serve as soft selectors to dynamically identify the most relevant information for the current input. This strategy can effectively mitigate the limitations imposed by context window sizes and enhance the model’s ability to precisely capture editing information.
@@ 3.1 Edit Compression
We adopt the concept of gisting [30], which is originally developed to compress input prompts into the representations of an extra, specially inserted token (the gist token). The condensed gist activation serves the same function as the original prompt and can be cached for later reuse, thereby improving computational and memory efficiency. While the original work primarily focuses on instruction tuning, we extend this idea to edit compression.
@@ 3.3 Training
In addition to token reweighting, we also adopt knowledge distillation [14] to transfer the teacher model’s knowledge about the edit information into the target model. Specifically, we apply the KL divergence to align the output distributions of the gist-contextualized student model with those of the edit-contextualized teacher model:
@@
Here, $g_{1},\dots,g_{n}$ denote the cached gists for all the edits. We apply the token reweighting only to the vanilla cross-entropy term in our final loss, since we found that it would degrade effective learning of “attend-to-nothing” behavior if combined with the KL part. The explicit training details are provided in Appendix B.
@@ 4.1 Experiment setting
To verify the effectiveness of our method in complex editing scenarios, we conduct experiments on five popular datasets in model editing: the dataset for multi-hop editing MQuAKE [49], the natural language editing dataset DUNE [1], the extended version of ZsRE [43, 46], which adds a portability test set to the original ZsRE [20], and the dataset containing ripple effect samples $\text{WikiData}_{counterfact}$ [2, 46]. We report edit success rate and portability for the extend-ZsRE and $\text{WikiData}_{counterfact}$, the results for 2, 3, and 4 edits for MQuAKE [49] and new information, scientific reasoning, and debiasing for DUNE [1]. More details about the datasets and evaluation metrics can be found in the Appendix C.1 and Appendix C.2, respectively.
@@
For baselines, we select representative methods demonstrated to be powerful in relevant surveys [43, 46]. For methods that directly edit the model weights, we include ROME [24], R-ROME [10], and MEMIT [25]; for methods that adopt explicit external memory, we include SERAC [29], and IKE [48]; for methods that train additional meta-model or use implicit external memory (stores activations or nerons, etc), we adopt MEND [28], GRACE [12] and KN [3]. We also include the traditional but powerful method, like fine-tuning, LoRA [16], and ICL, which directly concatenates all the edits as the prefix context. While some similarities exist between our method and RAG, they vary considerably in problem setting and methodology. A detailed analysis is given in Appendix D.1. We choose two representative open-source models for evaluation: Llama-3.2-1B  [^4] https://huggingface.co/meta-llama/Llama-3.2-1B and Qwen2.5-7B [42]. Unless otherwise specified, we adopt an edit batch size of 100 for batch editing. More details on the baseline implementation can be found in the Appendix C.3.
@@ 4.2 Main results
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits 2-edits 3-edits 4-edits Base Llama-3.2-1B 41.79 43.51 31.58 41.79 43.51 31.58 FT-M 55.32 56.59 52.22 51.89 50.15 44.1 LoRA [16] 67.63 68.84 60.21 50.29 47.88 47.15 ROME [24] 1.97 7.77 0.55 - - - R-ROME [10] 2.72 4.29 8.53 - - - MEMIT [25] 40.21 39.54 22.34 29.71 28.87 15.7 EMMET [11] 5.4 6.61 4.74 11.49 17.16 15.89 GRACE [12] 8.45 7.13 5.24 2.03 2.68 2.06 SERAC [29] 41.84 43.51 31.58 41.84 43.45 31.74 MEND [28] 39.88 35.23 33.31 35.55 30.29 29.89 ICL 59.23 59.00 51.63 50.06 49.86 42.37 InComeS $\textbf{71.19}_{20.19\%\uparrow}$ $\textbf{72.17}_{22.32\%\uparrow}$ $\textbf{72.62}_{40.65\%\uparrow}$ $\textbf{53.93}_{7.73\%\uparrow}$ $\textbf{52.79}_{5.88\%\uparrow}$ $\textbf{52.73}_{24.45\%\uparrow}$ Base Qwen2.5-7B 44.08 44.14 30.62 44.08 44.14 30.62 FT-M 69.89 74.88 74.96 50.31 49.04 50.39 LoRA [16] 36.95 34.28 29.02 16.41 24.22 23.12 ROME [24] 9.67 7.33 8.93 - - - R-ROME [10] 10.73 6.68 3.62 - - - MEMIT [25] 44.14 46.05 32.15 43.94 46.10 30.55 EMMET [11] 26.10 38.38 26.36 40.83 45.02 33.89 GRACE [12] 15.78 13.45 13.23 5.43 7.88 4.23 SERAC [29] 55.56 59.23 53.67 42.34 40.33 39.39 MEND [28] 34.23 45.34 30.25 39.88 35.45 34.21 ICL 69.76 76.91 74.54 53.53 50.54 44.77 InComeS $66.46_{4.73\%\downarrow}$ $71.24_{7.37\%\downarrow}$ $\textbf{76.54}_{2.68\%\uparrow}$ $\textbf{55.13}_{3.00\%\uparrow}$ $\textbf{53.48}_{5.82\%\uparrow}$ $\textbf{47.91}_{7.01\%\uparrow}$
Table 1: Results on MQuAKE [49]. The difference between InComeS and ICL is marked.
@@
One of our method’s advantages is its flexibility in handling a variety of editing contexts with different formats. Unlike many traditional editing methods like ROME [24] and MEMIT [25], which require the input to follow the triplet-like fact statement format, InComeS can take edits in free-text forms without explicitly labeled subjects and objects. To verify our method’s capability for such scenarios, we adopt the DUNE dataset, which includes natural-language form edits, and the results are shown in Table 2. Following the original paper of DUNE [1], we include fine-tuning, LoRA, SERAC [29], and ICL as our baselines. The result confirms our method’s capability to handle natural language edits. Interestingly, the raw model itself is a strong baseline in the batch editing scenario, which may demonstrate the fast-evolving model capabilities over the years.
Method Model Single Editing Batch Editing New info Scientific R. Debiasing New info Scientific R. Debiasing Base Llama-3.2-1B 56.85 55.87 32.73 56.85 55.87 32.73 FT-M 57.43 53.34 35.73 57.07 53.45 33.43 LoRA 53.65 50.87 36.73 56.77 54.66 35.83 SERAC [29] 52.67 48.96 46.73 50.76 47.32 36.52 ICL 58.67 55.84 56.65 56.94 55.68 39.46 InComeS $\textbf{60.00}_{2.27\%\uparrow}$ $\textbf{58.17}_{4.17\%\uparrow}$ $54.61_{3.60\%\downarrow}$ $\textbf{57.76}_{1.44\%\uparrow}$ $\textbf{56.46}_{1.40\%\uparrow}$ $\textbf{46.14}_{16.93\%\uparrow}$ Base Qwen2.5-7B 63.44 66.03 36.95 63.44 66.03 36.95 FT-M 64.83 67.77 48.73 64.27 64.56 41.51 LoRA 63.85 63.22 39.56 62.58 65.33 43.73 SERAC [29] 64.45 63.57 33.38 56.78 58.97 31.24 ICL 65.81 65.34 35.25 66.29 66.24 40.73 InComeS $\textbf{66.83}_{1.55\%\uparrow}$ $\textbf{68.02}_{4.10\%\uparrow}$ $\textbf{62.59}_{77.56\%\uparrow}$ $65.61_{1.03\%\downarrow}$ $\textbf{67.82}_{2.39\%\uparrow}$ $\textbf{56.69}_{39.18\%\uparrow}$
Table 2: Results on DUNE [1]. The difference between InComeS and ICL is marked.
@@
We further evaluate our method on two popular editing datasets that require reasoning abilities: $Wiki_{counterfact}$ [2, 46] and the extended ZsRE [43, 46]. Table 3 shows the results. [^5] We only present methods deemed representative and powerful on these two datasets, detailed version can be found in Table 7. Our primary focus is on portability, as it serves as the most representative metric for assessing a model’s comprehensive understanding of the editing information. Overall, our method achieves performance comparable to ICL and consistently outperforms other baselines. As expected, traditional editing methods such as fine-tuning and ROME exhibit high edit success rates; however, their portability scores generally lag behind the top-performing methods, highlighting a common limitation of these approaches. In contrast, ICL-based approaches that leverage the in-context learning capabilities of LLMs demonstrate superior performance in complex editing scenarios that require reasoning, owing to the enhanced context understanding of LLMs.
Method Model $\text{WikiData}_{counterfact}$ ZsRE-extended Edit Success Portability Edit Success Portability Base Llama-3.2-1B 21.28 / 21.28 19.73 / 19.73 30.06 / 30.06 40.17 / 40.17 FT-M 97.02 / 94.58 53.43 / 47.51 99.81 / 95.94 62.80 / 54.84 LoRA [16] 98.91 / 82.61 52.87 / 43.84 99.86 / 93.18 57.43 / 44.85 ROME [24] 94.33 / 44-33 40.44 / 44-33 95.41 / 44-33 46.04 / 44-33 MEMIT [25] 44-33 / 66.94 44-33 / 23.51 44-33 / 58.79 44-33 / 25.68 MEND [28] 44-33 / 26.66 44-33 / 21.06 34-33 / 43.33 34-33 / 30.77 GRACE [12] 33.27 / 25.06 14.33 / 10.51 32.00 / 24.44 12.73 / 10.91 IKE [48] 61.70 / 34-33 45.55 / 34-33 59.15 / 43-33 57.39 / 43-33 SERAC [29] 89.56 / 78.32 60.56 / 40.45 92.69 / 89.61 66.59 / 51.61 ICL 93.31 / 82.95 65.81 / 49.75 68.86 / 60.84 62.19 / 55.58 InComeS 91.16 / 76.81 65.15 / 45.66 97.22 / 87.09 70.70 / 52.23 Base Qwen2.5-7B 22.35 / 22.35 21.46 / 21.46 36.21 / 36.21 43.86 / 43.86 FT-M 98.93 / 90.18 49.39 / 43.13 99.51 / 92.60 50.04 / 46.41 LoRA [16] 77.31 / 72.22 37.04 / 31.65 86.88 / 77.78 28.61 / 24.13 ROME [24] 92.69 / 34-33 40.25 / 34-33 97.86 / 34-33 50.43 / 34-33 MEMIT [25] 34-33 / 91.16 34-33 / 39.85 34-33 / 93.28 34-33 / 49.97 MEND [28] 34-33 / 35.13 34-33 / 15.29 34-33 / 50.91 34-33 / 38.83 GRACE [12] 31.34 / 33.77 25.60 / 18.55 33.27 / 26.79 14.35 / 11.25 IKE [48] 96.40 / 34-33 75.33 / 34-33 99.75 / 34-33 83.17 / 34-33 SERAC [29] 91.79 / 80.68 51.12 / 41.26 91.12 / 82.56 62.41 / 52.63 ICL 90.24 / 85.28 66.99 / 51.66 71.75 / 71.57 66.10 / 64.57 InComeS 90.96 / 71.44 66.69 / 47.93 97.95 / 91.29 75.63 / 61.22
Table 3: Results for $WikiData_{counterfact}$ [2, 46] and ZsRE-extended [43, 46]. The data format of each cell is in "single/batch editing results". "-" means the methods are not designed for the corresponding settings. The best two statistics are marked.
@@
We further provide a scaling-up analysis to illustrate our method’s ability to generalize to larger numbers of edits, which is the main motivation of our modification over the ICL baseline. For this analysis, we use the COUNTERFACT dataset [25], as it provides a sufficient number of editing instances. We vary the number of edits from 100 to 1000, resulting in total token counts ranging from approximately 1.2K to 12K. The results are shown in Figure 3, which shows that InComeS consistently outperforms ICL, though the base models have already been pretrained over long contexts [42]. This finding suggests that the vanilla attention mechanism alone is insufficient to effectively comprehend and precisely select the required information from the context in complex editing scenarios. In contrast, our method demonstrates greater potential for handling large-scale edits through the unified compression and selection mechanism.